OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 7, -11, and 2 + 8i

OpenStudy (anonymous):

it is clear that two factors are \((x-7)\) and \((x+11)\)?

OpenStudy (anonymous):

the hard part is finding a polynomial with zeros at \(2+8i\) and its conjugate \(2-8i\) there are a couple ways to do it one is hard one is easy one is real real easy you choose

OpenStudy (anonymous):

real easy

OpenStudy (anonymous):

f(x) = x4 - 9x3 - 56x2 + 290x - 5236 f(x) = x4 - 9x3 + 56x2 - 290x + 5236 f(x) = x4 - 145x2 + 580x - 5236 f(x) = x4 - 25x2 + 580x - 5236